OpenStudy (anonymous):

Solve for x. 8 – 3x – x = 12

OpenStudy (anonymous):

Keep the like terms together. \[-3x-x = 12 - 8\] \[-4x = 4\] \[x = \frac{ 4 }{ -4 }\]

OpenStudy (aripotta):

or x = -1